OpenStudy (anonymous):

What is the equivalent expression of w^3x^-5y^4z^-2

OpenStudy (somethingawesome):

Negative exponents mean division, so x^-5 = 1/(x^5). Then \[ w^3 x^{-5} y^4 z^{-2} = \frac{w^3 y^4}{x^5 z^2}\]